OpenStudy (anonymous):

Find the missing side of the right triangle with sides: h = 5, a = 3

OpenStudy (anonymous):

4

OpenStudy (anonymous):

3-4-5 triangle 4

OpenStudy (anonymous):

that was fase

OpenStudy (anonymous):

\[3^2+4^2=5^2\] for sure

OpenStudy (anonymous):

if you didn't recognize the 3-4-5 right triangle you would write \[3^2+b^2=5^2\] \[9+b^2=25\] \[b^2=16\] \[b=4\]